if the sides of a square are increased by 5 feet, the new area is 81 square feet. How long is each side of the original square?
Answer by stanbon(75887)

if the sides of a square are increased by 5 feet, the new area is 81 square feet. How long is each side of the original square?
----
Original side:: x ; Original area = x^2
-----
New side:: x+5 ; New area = (x+5)^2
-----
Equation:
(x+5)^2 = 81
----
x+5 = 9
----
x = 4 ft (length of each side of the original square)
--------------
Cheers,
Stan H.
--------------